AB INITIO STUDY OF ION-PAIR STATES OF THE IODINE MOLECULE

\underline{VADIM A. ALEKSEEV}, Institute of Physics, St.Petersburg State University, Uljanovskaja St.1, Peterhof, 198504 St.Petersburg, Russia.

Ion-pair states of the I_{2}molecule have been the subject of many experimentals studies and to date all 18 states correlating with I^{+}(^{3}P_{J=2,1,0}, 1D_{2}) + I^{−}(^{1}S_{0}) asymptotes are known from experiment. This contribution reports on ab initio study of the I_{2} molecule with an emphasis on the ion-pair states. Figure shows experimental and calculated potentials of the ion-pair states correlating with I^{+}(^{3}P_{2}) + I^{−}(^{1}S_{0}) asymptote (energy is relative to I (^{2}P_{3/2}) + I (^{2}P_{3/2}) asymptote).